OpenStudy (anonymous):

6xy(1/2x^2-1/3xy+1/6y^2)

OpenStudy (jhannybean):

\[\large 6xy\left(\frac{1}{2x^2}-\frac{1}{3xy}+\frac{1}{6y^2}\right)\] Is this your problem?